Gd3+ has seven unpaired electrons, find its spin only magnetic moment. 


1
Expert's answer
2021-04-06T04:35:02-0400

"Magnetic Moment =\sqrt (n(n+2)"


Magnetic Moment ="\sqrt(7(7+2)=\sqrt 63"


Magnetic Moment =7.9BM